# A Disproof of Large Language Model Consciousness: The Necessity of Continual Learning for Consciousness

### v1

@@ 6 Discussion
6.1 Defining a Progressive Research Program on Consciousness
@@
The literature around formal requirements theories of consciousness, which so far has focused on falsification [doerig2019unfolding] and non-triviality [kleiner2021falsification] and also “toy” models [albantakis2025utility], may seem at first a small subfield of the neuroscience of consciousness focused on highly-theoretical “armchair” arguments. For this reason, others have argued that formal or theory-based arguments should take a backseat to empirical research, and we should not put strong constraints on theories of consciousness. [usher2023philosophical]
@@
The results herein show the opposite conclusion. Formal requirements for theories of consciousness to be testable offer the glimpse of a rare progressive subfield of consciousness research. To put it bluntly: the past two decades have seen a number of novel theories of consciousness advanced, and empirical research has, so far, not ruled out a single theory of consciousness that I’m aware of. Nor has it answered any urgent questions we desire of consciousness research, like the issue of LLM consciousness.
@@
It seems likely that many will approach these results by debating the definition of falsification, such as pointing out that while a simpler Popperian definition is used here [popper2005logic], in comparison Lakatos offers a more complex form of falsification [lakatos2014falsification]. Even if some more complicated conception survives the substitution argument (which we have no reason to think it should, given that it is just “refined falsification”), it is a move of last resort: if we have clear and simple falsification criteria for a science of consciousness, and we know certain theories do pass these without cases of pathological mismatches for identical input/output systems, and others do not, we should be very careful about relaxing the standards of science until it’s shown to be impossible to make progress without doing so.
@@
The point of this paper goes beyond its contents: the goal is to illustrate how a progressive approach to a science consciousness might take place via painstakingly working out all the requirements theories of consciousness need to satisfy in order not to be a priori falsified and yet still remain non-trivial and scientifically informative. That this already rules out a great number of contemporary theories, gives us the structure of a strong disproof of LLM consciousness, and points us toward the capability of continual learning, is the kind of actually surprising information that indicates a progressive research program.
@@
Indeed, this allows us to define a precise target for theoretical consciousness research: a Minimal Non-trivial and Falsifiable Theory of Consciousness (MNFTC). The MNFTC is a theory of consciousness that does not succumb to any of the logical problems around falsifiability and testability that contemporary theories may all fall into. It seems that equating learning with consciousness in some way may be the clearest path to this, although defining such a theory is left for future research.
@@ 6.2 Constraints on Theories Rule out the Consciousness of Many Artificial Systems (But Not All)
  • The “lenient dependency” required for theories of consciousness to escape issues of falsification (or unfalsifiability) must be continually true and present in non-trivial and non-falsified theories.
@@
One potential limitation is simply that these claims are too strong, indicating that there may be flaws in the reasoning here (if only due to my poor instantiation of it). Additionally, there is a large amount of future research to be done sketching out more formally what this progressive research program would consist of, delineating its boundaries, and also evaluating potential theories that actually fit those constraints, and connecting this to other work [kleiner2025logic]. Therefore, it’s worth noting there may be other cases of lenient dependency that are more exotic. However, plasticity has a strong claim by being already reasonable to include in inferences about consciousness. Speculative alternatives include, say, observer effects in quantum mechanics being somehow immune to unfolding but also still (again, somehow) reasonable to include in inferences [hameroff2014consciousness].
@@ 6.3 Continual Learning as Central to Consciousness
It is not a coincidence that LLMs, despite their intelligence, still feel like “ghosts” [karpathy2025ghosts]. If consciousness is truly based on continual learning (or requires it for some further property), then it makes sense that the substitution of a frozen input/output echo of humans would be technically impressive but strangely fail in ways hard to specify or quantify.
@@
Currently, even the best contemporary LLMs struggle on long-term context and agential tasks, like playing Pokémon without a hand-holding scaffold [angliss2025benchmark] Despite all their impressive gains on benchmarks and in real-world usage, they are still extremely inefficient in their learning (which requires implementing improvements in steps) and this does have significant functional effects [bell2025future]. Therefore, this has made many speculate about the necessity of some unknown type of continual learning as being necessary for artificial general intelligence [yildiz2024investigating].
Despite LLMs obvious intelligence on benchmarks and other tasks, it is still indisputable that even a human baby is orders of magnitude more data-efficient than an LLM [warstadt2023findings, frank2023bridging]. If LLMs are indeed non-conscious, this should tilt us toward theories of consciousness based on learning, since even animals are arguably much more data-efficient than contemporary AI.
@@
Of course, it is worth noting that there are some theories of consciousness that are already focused on learning and plasticity [cleeremans2011radical, o2025delay]. The results herein should make these prime candidates for exploration, adoption, and testing. It is a matter of future research to what degree these explicitly fit the requirements of continual lenient dependency, as some may be based only on past learning, rather than current learning. It is also worth noting in this context of continual learning that, intriguingly, the disproof does not rule out LLM consciousness during their training in the same way it does for the deployed “dead” artificial neural network; however, nor does it guarantee this in any way (i.e., training might not be similar to actual the kind of continual learning involved in consciousness), and so this remains an area reserved for future research.
@@
Overall, a view of the biological brain as being a gestalt continuously (and significantly) updated by every experience is, at least, scientifically evocative, and not contradicted by the neuroscience of learning. E.g., place fields can develop in a single trial [bittner2017behavioral], and significant plasticity can occur over seconds [chaloner2022multiple]; even the Overfitted Brain Hypothesis hypothesizes that dreams may have evolved to combat daily overfitting [hoel2021overfitted].
@@
Therefore, it may well be that from the beginning evolution has been focused on continual learning in an organism’s appropriate domain (its umwelt), rather than intelligence. This would explain our current uncertainty when high intelligence—at least on things like tests and benchmarks—does not automatically lead to consciousness.

### v2

@@ 6 Discussion
6.1 A Progressive Research Program on Consciousness
@@
The literature around formal requirements theories of consciousness, which so far has focused on falsification [31] and non-triviality [45] and also “toy” models [4], may to outsiders at first seem a small subfield of the neuroscience of consciousness focused on highly-theoretical “armchair” arguments. For this reason, some in the literature have argued that formal or theory-based arguments should take a backseat to empirical research, and we should not put strong constraints on theories of consciousness. [72]
@@
The results herein show the opposite. Formal requirements for theories of consciousness to be testable offer the glimpse of a rare progressive subfield of consciousness research. To put it bluntly: the past two decades have seen a large number of novel theories of consciousness advanced, and empirical research has, so far, not ruled out a single theory of consciousness that I’m aware of. Nor has it answered any urgent questions we desire of consciousness research, like the issue of LLM consciousness.
@@
It seems likely that some will argue against these results by debating the definition of falsification, such as pointing out that while a simpler Popperian definition is used herein [62], in comparison Lakatos offers a more complex form of falsification [55]. However, even if some more complicated conception of falsification complicates the Kleiner-Hoel dilemma (we have no reason to think it would rule it out, given that such alternatives are often just “refined” versions of Popperian forms of falsification), it would still be a move of last resort: if we have clear and simple falsification criteria for a science of consciousness, and we know certain theories do pass these without cases of pathological mismatches for identical input/output systems, and others do not, we should be very careful about relaxing the standards of science until it’s shown to be impossible to make progress without doing so.
@@
The point of this paper goes beyond its contents: the goal is to illustrate how a progressive approach to a science of consciousness might take place via painstakingly working out all the requirements theories of consciousness need to satisfy in order not to be a priori falsified and yet still remain non-trivial and scientifically informative. That this already rules out a great number of contemporary theories, gives us the structure of a strong disproof of LLM consciousness, and points us toward the capability of continual learning, is the kind of actually surprising information that indicates a progressive research program.
@@
Indeed, this allows us to define a precise target for theoretical consciousness research: a Minimal Non-trivial and Falsifiable Theory of Consciousness (MNFTC). An MNFTC is a theory of consciousness that, via satisfying continual lenient dependency, does not succumb to any of the logical problems around falsifiability and testability that the super-majority of contemporary theories fall into. It seems that somehow equating learning with consciousness in some way may be the clearest path to this, although defining such a theory is left for future research.
@@ 6.2 Constraints on Theories Rule out the Consciousness of Many Artificial Systems (But Not All)
  • The “lenient dependency” required for theories of consciousness to escape issues of falsification (or unfalsifiability) must be continually true and present in falsifiable and non-trivial theories.
@@
One potential limitation is simply that these claims are too strong, indicating that there may be flaws in the reasoning here (if only due to my poor instantiation of it). Additionally, there is a large amount of future research to be done sketching out more formally what this progressive research program would consist of, delineating its boundaries, and also evaluating potential theories that actually fit those constraints, and connecting this to other work [50]. Therefore, it’s worth noting there may be other cases of lenient dependency beyond continual learning that are more exotic (speculative alternatives include, say, observer effects in quantum mechanics being somehow immune to substitutions [36]). However, learning has a strong claim by being already reasonable to include in inferences about consciousness: could there really be experiences that leave no trace whatsoever?
@@ 6.3 Continual Learning as Central to Consciousness
It is not a coincidence that LLMs, despite their intelligence, still feel like “ghosts” [43]. If consciousness is truly based on continual learning (or requires it for some further property), then it makes sense that the substitution of a frozen input/output echo of humans would be technically impressive but fail as our substitutions in ways that are hard to specify or quantify. Often science proceeds spurred on by metaphors from technology: LLMs being close substitutions to humans for many input/output text-based or screen-based tasks, but then also being somehow incomplete substitutions in ways difficult to quantify or specify, seems to preface the results herein about how input/output equivalence is a less valid target than first appears.
@@
Currently, even the best contemporary LLMs struggle on long-term context and agential tasks, like playing Pokémon without a hand-holding scaffold [6]. Despite all their impressive gains on benchmarks and in real-world usage, they are still extremely inefficient in their learning and this does have significant functional effects [13]. This has led to many speculating about the necessity of some unknown type of continual learning as being necessary for true artificial general intelligence [74], as it is indisputable that even a human baby is orders of magnitude more data-efficient than an LLM [73, 33]. If LLMs are indeed non-conscious, this too should tilt us toward theories of consciousness based on learning, since even far less intelligent animals are much more data-efficient in their learning than contemporary AI.
@@
Of course, it is worth noting that there are some theories of consciousness that are already focused on learning and plasticity [25, 60]. The results herein should make these prime candidates for exploration, adoption, and testing. It is a matter of future research to what degree existing theories explicitly fit the requirements of continual lenient dependency, as some may be based only on past learning, rather than current learning.
Intriguingly, the disproof herein does not rule out LLM consciousness during their training in the same way it does for the deployed “dead” artificial neural network; however, nor does it guarantee their consciousness in any way (i.e., training might not be similar to the actual kind of continual learning involved in consciousness. nor be technically continual in practice), and so this remains an area reserved for future research. When it comes to training, it should be noted that there are detectable differences between how humans vs. LLMs learn, already indicating that the learning process itself may differ significantly and indicating that LLM consciousness may not hold during training [61].
@@
Finally, with regard to humans, a view of the biological brain as being a gestalt continuously updated by every experience is, at least, scientifically evocative, and not contradicted by the neuroscience of learning. E.g., place fields can develop in a single trial [15], and significant plasticity can occur over seconds [24]; even the Overfitted Brain Hypothesis hypothesizes that dreams may have evolved to combat daily overfitting [37]. It may indeed be that every experience must leave a trace for it to be an experience at all.
@@
Therefore, it may well be that, from the beginning, evolution has been focused on continual learning in an organism’s appropriate domain (its umwelt), rather than intelligence. This would explain our current uncertainty when high intelligence—at least on things like tests and benchmarks—does not automatically lead to consciousness.